a force of 72 n acting upon a given object results in an acceleration of 8 m/s squared.
if the magnitude of the force acting upon the same object is changed to 63 n, what will the objects acceleration become?

would it be 7 m/s squared

Answer by dabanfield(803)

a force of 72 n acting upon a given object results in an acceleration of 8 m/s squared.
if the magnitude of the force acting upon the same object is changed to 63 n, what will the objects acceleration become?
would it be 7 m/s squared
Force = mass * acceleration (F = m*a)
Let m be the mass of the object.
We know that 72 = m*8 so m = 72/8 = 9.
If the acceleration is 63 then we have:
63 = m*a
63 = 9*a
a = 63/9 = 7 m/s sq